# Beyond Labeling Oracles: What does it mean to steal ML models?

### v1

@@ 6 Conclusion
In this paper, we show that ME attacks can often be more expensive for the adversary than developing a model from scratch. We demonstrate that the performance of an attacker with a reasonable query budget is bounded by their access to IND data. Yet, given sufficient IND data, the victim model mainly serves as a labeling oracle, which in turn is not necessarily more cost-effective than online labeling services. Augmenting the IND data with data from another distribution relies on informative responses for task-irrelevant queries. We show that decorrelation of OOD from IND responses changes the attacker-victim dynamic, where attacks become much less cost-effective, forcing the attacker to collect IND data. Although ME attacks represent a serious threat in practice, our work highlights that corresponding risks can be reduced by making extraction expensive with an appropriate incentive mechanism design. We leave this implementation for future work.
@@ Acknowledgments
Appendix A Primer on ML costing
Current model stealing literature never formally defines what makes a model extraction attack successful. While for high-fidelity extraction attacks an attacker explicitly expects to extract a model within a given error, accuracy-driven model extraction can be considered successful if stealing a model costs less than developing it.
The overall cost of producing and deploying a model can be broken down into six main parts:
  1. 1
    Data and Model storage
  2. 2
    Training and Deployment hardware
  3. 3
    Data collection
  4. 4
    Data labeling
  5. 5
    Model training
  6. 6
    Data and Model management
Out of the six, current model extraction really only focuses on data collection, labeling and training; training costs are usually only mentioned in passing and the vast majority of papers scale their attacks with just data collection. Although rationality of management and hardware costs in model stealing are indeed interesting, they fall outside of the current threat models in the literature.
Now we turn to CIFAR10 benchmark as a costing case study. Here we assume CIFAR10 dataset with 60k samples that can be labeled for $35\$*50+10*25\$=2000\$$ with Google Cloud  [^1] https://cloud.google.com/ai-platform/data-labeling/pricing and $0.04\$*60000=2400\$$ with Amazon Sagemaker  [^2] https://aws.amazon.com/sagemaker/data-labeling/pricing/. Data collection here is practically free, since per category one can scrape the internet freely.
We model the defender cost as $c=n*(pl+cc)$ and the attacker cost as $c_{a}=n_{a}*(pl_{a}+cc_{a})$, where $n$ is the number of points to annotate, $pl$ is the per-label cost and $cc$ is the data collection cost. We assume that the attacker wants to succeed in the attack and thus attacker’s cost is less than that of a defender $n*(pl+cc)>n_{a}*(pl_{a}+cc_{a})$.
First, we consider the attacker who extracts a model with DFME with 20 million queries. If we assume that the defender used Google Cloud to label their data, attacker’s labels have to cost less than $0.00012\$$:
$20M*pl_{a}<2,400\$\rightarrow pl_{a}<0.00012\$$ (1)
The queries in DFME are synthetic, therefore there is no data collection cost ($cc_{a}=0$). Next, we consider a model extraction attack that only utilizes $5\%=3000$ datapoints of prior knowledge, and no additional queries of any sort. Here, attacker breaks even at $0.8\$$ per-label. Note that with 5% of data it is indeed possible to extract a model, while remaining well within a reasonable budget in the current labeling market. At the same time, extraction with DFME stops being cost-effective when the attacker queries are even minimally priced by the defender.
This case study suggests the need for pricing of attacker queries and that model extraction might not be the most cost-effective solution, even for simple, commonly used benchmarks such as CIFAR10.
It is worth noting that this relationship does not always hold true in other cases. Medical data, for example, requires significant upfront equipment investment and needs prolonged interactions with data suppliers. This in turn means that the costs will be rather dominated by the data acquisition, as opposed to the labeling and such case will exhibit a different behaviour.
Appendix B A note on current literature
@@
Appendix C Effect of surrogate dataset over attack performance
@@
Appendix D Attack convergence rate comparison
@@
Appendix E Sampling complexity intuition
@@ E.1 Sampler bias toy example
E.2 Sampling complexity in real models
In the previous section, we explore the sampling complexity of an attacker in a toy setting. In this section, we extend this notion to a more realistic scenario of estimating the sampling complexity for our real victim models. For this, we attempt to estimate the in-distribution volume and the complexity of sampling within it by measuring the number of random queries in this volume. Similarly to our OOD component, described in detail in Section 5.3.1 and Appendix F, we define a query as in-distribution, i.e. inside the volume, by observing the model’s confidence in predicting this query. High confidence queries, above some predefined confidence threshold, are considered to be in-distribution. Therefore, for the task of volume estimation, we measure the percentage of random queries that are above this threshold. In this section, we use a threshold value of $90$.
@@ E.2 Sampling complexity in real models
Appendix F OOD full implementation details
In this section, we elaborate on the implementation details behind our OOD component, described shortly in Section 5.3.1. As described before, given the original victim model $\mathcal{V}_{o}$, we create a hybrid victim model $\mathcal{V}_{h}$ by combining the original victim model $\mathcal{V}_{o}$ with an additional module with different, or additional, decision boundaries $\mathcal{V}_{f}$. For each query $x$, we apply a decision rule $R$ over $x$, to determine which of the two modules, $\mathcal{V}_{o}$ or $\mathcal{V}_{f}$, should be used for predicting this query.
The decision rule $R$ is implemented by applying a threshold value $\tau$ over the prediction confidence of $\mathcal{V}_{o}$. If a query $x$ has a low prediction confidence, we define it as an OOD query and return the prediction $\mathcal{V}_{f}(x)$. Otherwise, we define it as IND and return $\mathcal{V}_{o}(x)$. We calibrate $R$ by increasing the model’s Softmax temperature to 2, as discussed in more detail in Section F.2. As mentioned in  Section 5.3.1, thresholding the prediction confidence serves as a naive approximation for OOD detection, and using more sophisticated OOD detection methods will only increase the effect of our OOD component. As such, the exploration of different decision rules is out of scope for this work.
@@ Appendix F OOD full implementation details
The construction of our OOD component requires the evaluation of a subset of the training data used for fitting the GMMs. The evaluation, fitting, and clustering process are done only once; hence, it is relatively computationally inexpensive. During inference, each query directed to the OOD component adds some computational cost of computing its feature representation, and performing a nearest neighbor search. However, this is only true for OOD samples and false-positive IND samples. Most legitimate users’ queries, i.e. the true-positive queries, are completely unaffected by our OOD component.
@@ F.1 Effect of anchor points and permutations
1 anchor no perm 1 anchor 1 perm 5 anchors 1 perm 5 anchor 5 perms
SVHN
Refer to caption Refer to caption Refer to caption Refer to caption
DFME
Refer to caption Refer to caption Refer to caption Refer to caption
random (MNLI)
Refer to caption Refer to caption Refer to caption Refer to caption
nonsensical (MNLI)
Refer to caption Refer to caption Refer to caption Refer to caption
wiki (MNLI)
Refer to caption Refer to caption Refer to caption Refer to caption
F.2 Softmax temperature influence
In order to calibrate $R$ to better distinguish between IND and OOD samples, we increase the model’s Softmax temperature to 2. By doing so, we force the model to be less confident, which results in more queries being predicted by $\mathcal{V}_{f}$ and not $\mathcal{V}_{o}$. As shown in Figure 9, a temperature value of 2 indeed detects most of the OOD queries. In Figure 11 we demonstrate the effect of the temperature value over the behaviour of our OOD component, and show that a better calibrated OOD component indeed further lowers the benefit the attacker gains by utilizing OOD queries.
@@ F.2 Softmax temperature influence
Figure 11: As our OOD component is distinguishing between in and out of distribution queries based on the model confidence, we evaluate the effect of decreasing the model’s confidence by increasing the softmax temperature T from 1 to 2. Results shown in this plot show that this indeed even further lowers the benefit the attacker gains by utilizing OOD queries.
@@ F.3 Impact of $\tau$
The value of the confidence threshold $\tau$ determines the utility – extraction difficulty trade-off. In Section 5.3.2 we evaluate our OOD component on different values of $\tau$, and observe the effect each has over the attack performance. In order to verify that our OOD component is not trivially rejecting all inputs, we additionally measure the false-positive rate (FPR) for the different threshold values. In Table 1, we detail the relation between the threshold value, the FPR, and the effect on the victim’s test accuracy. For lower $\tau$ values, most queries are predicted using the original $\mathcal{V}_{o}$, i.e. the victim model’s utility is barely harmed by adding the OOD component since most queries are answered by the real model. In this case, the attacker’s performance is barely affected, as it can still get informative responses by issuing OOD queries. As $\tau$ increases, more queries– including some IND ones– are predicted by the “fake” model $\mathcal{V}_{f}$. This makes it more difficult for an attacker to infer which decision boundaries are IND. We discussed the underlying complexity in detail in Sections 5.1 and 5.2. As can be seen, the attack accuracy dramatically decreased, which verifies that the OOD behaviour leaks almost no information about the IND behaviour. To further demonstrate this, we present in Figure 12 a comparison between the labels predicted by $\mathcal{V}_{o}$ and those predicted by $\mathcal{V}_{f}$ for actual OOD queries (i.e. true positives) as well as tail-of-distribution IND samples that were detected as OOD (i.e. false positives). We show this comparison for an attacker that utilizes additional SVHN queries, $30\%$ prior knowledge, and for a threshold of $\tau=95$. It is clear that, although the “fake” labels are heavily biased towards one class for the additional queries, in both cases, they are uncorrelated with victim models’ predicted labels.
@@ Appendix G Knockoff Nets Evaluation
G.1 Baseline Attack
We begin by evaluating our baseline attacker, following the setting described in Section 4.2. We assume the attacker has access to either a randomly or adaptively sampled subset of the victim’s true training dataset. The results, presented in Figure 13, align with our previous findings and show that in most cases, an attacker with access to more than $50\%$ can almost fully extract the model. These results align with the results presented in Orekondy et al.’s Figure 5.
As in Section 4.2, we compare the attack performance using soft-label access to the victim model to that of an attacker with access to the real (ground truth) labels or label-only access to the victim model. This comparison provides the same evidence as described in Section 4.2. It demonstrates that attacking the victim model is merely using the victim model as a labeling oracle in the absence of access to the real labels, and the attacker does not gain much additional benefit. This phenomena was also observed by Orekondy et al..
@@ G.1 Baseline Attack
G.2 Surrogate dataset
We investigate the attack performance gain from the use of additional queries, sampled from a data distribution that differs from the train distribution. Following Orekondy et al. we use ImageNet [Deng et al., 2009] as the surrogate dataset, from which we sample randomly or using the adaptive strategy. We follow the setting described in Section 4.3, and fix the query budget to be the size of the original training dataset. The results, presented in Figure 14 align with the results in Figure 3, as they show that as the attacker has more prior knowledge, it benefits less from utilizing additional queries.
We do observe that in the lower prior knowledge settings, the ImageNet additional queries do provide a significant improvement, especially for the Indoor67 and Caltech256 datasets. We hypothesize that this is due to a high similarity between the distributions. The results clearly demonstrate that the attacker’s performance is dominated by the access to IND data. The effect of the query selection strategy is inferior to that of data access, which aligns with the main findings of our paper.
(a) Indoor67
(b) CUBS200
(c) Caltech256
G.3 Limiting Out-Of-Distribution Effectiveness
We explore the reliance of the Knockoff Nets attacker on the implicit assumption that the IND decision boundaries can be inferred from the OOD ones. For this, we limit the informativeness of the OOD ImageNet queries by incorporating our OOD detection component. We follow the same setting as in Section 5.3.2 and evaluate the effect using different threshold values $\tau$. We set the model’s softmax temperature to $2$, and use $\|D_{train}\|$ additional queries.
Figure 15 demonstrates that limiting the informativeness of OOD queries indeed decreases the benefit the attacker gained by utilizing additional ImageNet queries, even to the point of performing worse than the baseline in the higher prior knowledge settings.
(a) Indoor67 (random)
(b) CUBS200 (random)
(c) Caltech256 (random)
(d) Indoor67 (adaptive)
(e) CUBS200 (adaptive)
(f) Caltech256 (adaptive)
Figure 15: The effect of applying our OOD component with different values of $\tau$ against an attacker that utilizes $\|D_{train}\|$ additional ImageNet queries. When comparing the results to the original setting (real model), where the OOD region is unmodified, we can see a clear decrease in the attack accuracy.
Appendix H NLP distribution similarity
The results presented in Figure 4 show that the effect of the OOD component is weaker for some of the settings in our NLP task, specifically the nonsensical and wiki queries. This might seem confusing. The wiki queries are broadly equivalent to the surrogate dataset queries that are commonly used in the vision domain, however they exhibit a different response to our OOD component. As both random images and random nonsensical sentences have no real meaning, one might expect similar behaviour between them. In the past, nonsensical queries were even used as a benchmark under the name of random queries e.g. by Truong et al..
This difference between the vision and NLP domains is due to the similarity between the true data distribution and the aforementioned query distribution. Although the queries are drawn from either a random nonsensical distribution in the nonsensical case, or a from a different corpus in the wiki case, we observe that many common words are shared between all three distributions. This results in similar model behaviour between the OOD queries and the true data. It can explain the success of the queries in this domain, which comes in significant contrast to the lack of success of random or some surrogate queries in the vision domain, where the input space is a continuous pixel space rather than a (relatively small) discrete dictionary of words. The similarity between distributions results in high confidence values for both in-distribution and out-of-distribution queries, making it difficult to separate them and apply our OOD component. We show the confidence distribution for each query type in Figure 16. In the case of the random queries, where each letter is sampled, and the sentences are not composed of real words, we can observe a significantly lower gain from the additional queries, which is more in line with the findings from the vision domain.

### v3

@@ 6 Conclusion
In this paper we investigate the common assumption that ME attacks are more cost-effective than training a model from scratch. Primarily, it is assumed that OOD data can be utilized for reducing data collection costs, and by using a limited number of queries to the victim model both labeling costs and training costs can be reduced as well. We show that often this is not the case. We demonstrate that the performance of an attacker with a reasonable query budget is bounded by their access to IND data. Augmenting the IND data with data from another distribution relies on informative responses for task-irrelevant queries. We show that decorrelation of OOD from IND responses changes the attacker-victim dynamic, where attacks become much less cost-effective, forcing the attacker to collect IND data. We show that given sufficient IND data, the victim model mainly serves as a labeling oracle, which in turn is not necessarily more cost-effective than online labeling services. We thus conclude that the adversarial goals and incentives of ME attacks should be redefined.
@@ Acknowledgments
Appendix A A note on current literature
@@
Appendix B Effect of surrogate dataset over attack performance
Appendix C Attack convergence rate comparison
@@
Appendix D Effect of response informativeness
In Section 5.5 we discussed the role of ME in reducing data labeling costs. We found that, in the setting considered in this work, ME attacks serve as a labeling oracle. To demonstrate this, we showed that when training the baseline attacker with soft-label access to the victim model, i.e,. when the victim model responds to each issued query with a full-probability vector, we did not get any significant improvement over cases where the attacker was trained with label-only access or even the real ground truth labels. In order to further emphasize the limited possible gain an attacker can get by obtaining labeling through a ME attack, we compare the difference in attack performance between soft-label access and label-only access for attackers that can utilize additional queries from other query distribution.
Similar to the results presented in Figure 2, we evaluate this for a constant query budget, fixed to the size of the original training set. Due to the high computational cost of DFME, we omit this setting in this evaluation.
The results, presented in Figure 9, demonstrate that soft-labels aid the attacker mainly in the lower prior knowledge settings and for query distributions that are significantly different from the true distribution. In other cases, the victim model serves as a labeling oracle, and the soft labels provide little gain over the label-only access. For example, in the CIFAR-10 case, an attacker that utilizes additional random queries and has access to $1\%$ of prior knowledge can improve it’s performance by additional $20.8\%$, from $26.43\%$ to $47.24\%$ attack accuracy. However, the same attacker with $10\%$ prior knowledge can only improve by $1.25\%$, from $81.31\%$ to $82.56\%$, and another attacker with $1\%$ prior knowledge that can utilize SVHN queries can only increase by $7.19\%$, from $68.14\%$ to $75.33\%$. In the case of MNLI, the benefit of soft labels is even weaker, due to the similarity between the distributions, as further discusses in Appendix H.
(a) CIFAR-10
(b) MNLI
Figure 9: Comparison between the attack performance in the soft-label setting, i.e., full probability vector, and the label-only setting. Evaluated for the baseline attacker, that utilizes only prior knowledge, and attackers that can utilize additional queries from other distributions (random images, SVHN, and CIFAR-100 in the case of CIFAR-10, and random sentences, nonsensical sentences, and wiki for the MNLI setting). Results demonstrate that soft-labels aid the attacker mainly in the lower prior knowledge settings and for query distributions that are significantly different from the true distribution. In other cases, the victim model serves as a labeling oracle, and the soft labels provide little gain over the label-only access.
@@ E.1 Sampler bias toy example
E.2 Sampling complexity in real models
In the previous section, we explore the sampling complexity of an attacker in a toy setting. In this section, we extend this notion to a more realistic scenario of estimating the sampling complexity for our real victim models. For this, we attempt to estimate the in-distribution volume and the complexity of sampling within it by measuring the number of random queries in this volume. Similarly to our OOD control module, described in detail in Appendix F, we define a query as in-distribution, i.e. inside the volume, by observing the model’s confidence in predicting this query. High confidence queries, above some predefined confidence threshold, are considered to be in-distribution. Therefore, for the task of volume estimation, we measure the percentage of random queries that are above this threshold. In this section, we use a threshold value of $90$.
Appendix F Controlling OOD informativeness - full implementation details
In this section, we elaborate on the implementation details behind our methodology for evaluating the effect of limiting the utility of OOD queries, described shortly in Section 4.3. Given the original victim model $\mathcal{V}_{o}$, we create a hybrid victim model $\mathcal{V}_{h}$ by combining the original victim model $\mathcal{V}_{o}$ with an additional module with different, or additional, decision boundaries $\mathcal{V}_{f}$. For each query $x$, we apply a decision rule $R$ over $x$, to determine which of the two modules, $\mathcal{V}_{o}$ or $\mathcal{V}_{f}$, should be used for predicting this query.
The decision rule $R$ is implemented by applying a threshold value $\tau$ over the prediction confidence of $\mathcal{V}_{o}$. If a query $x$ has a low prediction confidence, we define it as an OOD query and return the prediction $\mathcal{V}_{f}(x)$. Otherwise, we define it as IND and return $\mathcal{V}_{o}(x)$. We calibrate $R$ by increasing the model’s Softmax temperature to 2, as discussed in more detail in Section F.2. Thresholding the prediction confidence serves as a naive approximation for OOD detection, and using more sophisticated OOD detection methods will only increase the effect of modifying the OOD behaviour. As such, the exploration of different decision rules is out of scope for this work.
@@ Appendix F Controlling OOD informativeness - full implementation details
This construction requires the evaluation of a subset of the training data used for fitting the GMMs. The evaluation, fitting, and clustering process are done only once; hence, it is relatively computationally inexpensive. During inference, each query directed to the fake model $\mathcal{V}_{f}$ adds some computational cost of computing its feature representation, and performing a nearest neighbor search. However, this is only true for OOD samples and false-positive IND samples. Most legitimate users’ queries, i.e. the true-positive queries, are completely unaffected by the introduction of this additional module.
It is important to note that while our method decreases the utility of existing ME attacks, as shown in our results, we do not present it as a defense mechanism but rather as a method for evaluating our hypothesis. An efficient defense mechanism can be designed based on these principles, with some additional effort. However, this is out of scope for our paper. We propose a method for measuring the reliance of ME attacks on the informativeness of OOD regions, as part of a bigger discussion over the role of prior knowledge and the common methodology of ME attacks, and do not investigate the utilization of it as part of any defense mechanism.
@@ F.1 Effect of anchor points and permutations
Figure 12: As our OOD detection rule is distinguishing between in and out of distribution queries based on the model confidence, we evaluate the effect of decreasing the model’s confidence by increasing the softmax temperature T from 1 to 2. Results shown in this plot show that this indeed even further lowers the benefit the attacker gains by utilizing OOD queries.
F.2 Softmax temperature influence
In order to calibrate $R$ to better distinguish between IND and OOD samples, we increase the model’s Softmax temperature to 2. By doing so, we force the model to be less confident, which results in more queries being predicted by $\mathcal{V}_{f}$ and not $\mathcal{V}_{o}$. As shown in Figure 11, a temperature value of 2 indeed detects most of the OOD queries. In Figure 12 we demonstrate the effect of the temperature value over the behaviour of our controlled informativness experiment and show that a better calibrated decision rule indeed further lowers the benefit the attacker gains by utilizing OOD queries.
@@ F.3 Impact of $\tau$
The value of the confidence threshold $\tau$ determines the utility – extraction difficulty trade-off. In Section 4.3 we evaluate our OOD control mechanism on different values of $\tau$, and observe the effect each has over the attack performance. In order to verify that we are not trivially rejecting all inputs, we additionally measure the false-positive rate (FPR) for the different threshold values. In Table 1, we detail the relation between the threshold value, the FPR, and the effect on the victim’s test accuracy. For lower $\tau$ values, most queries are predicted using the original $\mathcal{V}_{o}$, i.e. the victim model’s utility is barely harmed since most queries are answered by the real model. In this case, the attacker’s performance is barely affected, as it can still get informative responses by issuing OOD queries. As $\tau$ increases, more queries– including some IND ones– are predicted by the “fake” model $\mathcal{V}_{f}$. This makes it more difficult for an attacker to infer which decision boundaries are IND. We discussed the underlying complexity in detail in Sections 4.1 and 4.2. As can be seen, the attack accuracy dramatically decreased, which verifies that the OOD behaviour leaks almost no information about the IND behaviour. To further demonstrate this, we present in Figure 13 a comparison between the labels predicted by $\mathcal{V}_{o}$ and those predicted by $\mathcal{V}_{f}$ for actual OOD queries (i.e. true positives) as well as tail-of-distribution IND samples that were detected as OOD (i.e. false positives). We show this comparison for an attacker that utilizes additional SVHN queries, $30\%$ prior knowledge, and for a threshold of $\tau=95$. It is clear that, although the “fake” labels are heavily biased towards one class for the additional queries, in both cases, they are uncorrelated with victim models’ predicted labels.
@@ Appendix G Knockoff Nets Evaluation
Similar to the evaluations in Section 5.1, we show in Figure 14 that ME indeed works, and the model can be approximated with blackbox query access. To better visualize this, we compare the performance of a baseline attacker, that only uses its prior knowledge over the distribution, with that of the OOD-based attacker. Here, the baseline attacker is assumed to have access to either a randomly or adaptively sampled subset of the victim’s true training dataset. Following Orekondy et al. we use ImageNet [Deng et al., 2009] as the surrogate (additional) dataset, from which we either sample randomly or using the adaptive strategy. We follow the setting described in Section 5.1, and fix the query budget to be the size of the original training dataset. The results, presented in Figure 14 align with the results in Section 5.1, as they show that as the attacker has more prior knowledge, it benefits less from utilizing additional queries. We do observe that in the lower prior knowledge settings, the ImageNet additional queries do provide a significant improvement, especially for the Indoor67 and Caltech256 datasets. We hypothesize that this is due to a high similarity between the distributions. This also aligns with the main findings of our paper.
(a) Indoor67 (random)
(b) CUBS200 (random)
(c) Caltech256 (random)
(d) Indoor67 (adaptive)
(e) CUBS200 (adaptive)
(f) Caltech256 (adaptive)
Figure 15: The effect of limiting the OOD informativeness with different values of $\tau$ against an attacker that utilizes $\|D_{train}\|$ additional ImageNet queries. When comparing the results to the original setting (real model), where the OOD region is unmodified, we can see a clear decrease in the attack accuracy.
Figure 14 additionally answers the questions described in Section 5.2 and Section 5.3. Namely, ME can be used with only a few queries, as the attack is succesfull both when bounding the query budget to the size of the original training set, and when using just a smaller fraction of IND samples. Moreover, the results also show that the use of ImageNet queries as OOD data is indeed effective and can help reduce the data costs involved with using expensive IND samples.
However, as we discussed in Section 4 and further evaluated in Section 5.4, the answers to the previous questions depends on the implicit assumption that the IND decision boundaries can be inferred from the OOD ones. We evaluate this assumption in this case study as well, following the methodology described in Section 4.3. We set the model’s softmax temperature to $2$, and use $\|D_{train}\|$ additional queries.
Figure 15 demonstrates that limiting the informativeness of OOD queries indeed decreases the benefit the attacker gained by utilizing additional ImageNet queries, even to the point of performing worse than the baseline in the higher prior knowledge settings. This again aligns with our findings in Section 5.4 and suggests that the attacker can not acheive both goals, and in the absence of prior knowedlge, corresponding to higher data collection costs, a high query budget must be used.
Aligning with the findings reported in Section 5.5, we present in Figure 16 a comparison between the attack performance using soft-label access to the victim model to that of an attacker with access to the real (ground truth) labels or label-only access to the victim model. This comparison provides the same evidence as described in Section 5.5. It demonstrates that attacking the victim model is merely using the victim model as a labeling oracle in the absence of access to the real labels, and the attacker does not gain much additional benefit. This phenomena was also observed by Orekondy et al..
@@
Appendix H NLP distribution similarity
The results presented in Figure 4 show that the effect of limiting the leakage from OOD behaviour to IND behaviour is weaker for some of the settings in our NLP task, specifically the nonsensical and wiki queries. This might seem confusing. The wiki queries are broadly equivalent to the surrogate dataset queries that are commonly used in the vision domain, however they exhibit a different response in our evaluation. As both random images and random nonsensical sentences have no real meaning, one might expect similar behaviour between them. In the past, nonsensical queries were even used as a benchmark under the name of random queries e.g. by Truong et al..
This difference between the vision and NLP domains is due to the similarity between the true data distribution and the aforementioned query distribution. Although the queries are drawn from either a random nonsensical distribution in the nonsensical case, or a from a different corpus in the wiki case, we observe that many common words are shared between all three distributions. This results in similar model behaviour between the OOD queries and the true data. It can explain the success of the queries in this domain, which comes in significant contrast to the lack of success of random or some surrogate queries in the vision domain, where the input space is a continuous pixel space rather than a (relatively small) discrete dictionary of words. The similarity between distributions results in high confidence values for both in-distribution and out-of-distribution queries, making it difficult to separate them and apply our method. We show the confidence distribution for each query type in Figure 17. In the case of the random queries, where each letter is sampled, and the sentences are not composed of real words, we can observe a significantly lower gain from the additional queries, which is more in line with the findings from the vision domain.
(a) SVHN
(b) SST-2
Figure 18: The baseline attacker is able to successfully extract some victim models very easily, with as little as $5\%$ of the training data. We hypothesize that this is due to the inherent linearity of these datasets.
Figure 19: We investigate whether the victim model size has any impact on the attack success. We observe no impact when increasing (ResNet-50-8x) or decreasing (ResNet-18-8x) the model size. All models are incredibly over-parameterized and, therefore, the differences between the models are insignificant in terms of the attack complexity.